# Cross-Encoder Rediscovers a Semantic Variant of BM25

### v2

@@ 5 Validation of BM25-like Computation
To validate whether Relevance Scoring Heads perform a BM25-style computation as hypothesized in §4.1, we first formalize the hypothesized function of the heads and the information flowing into them as a BM25-style linear function. Next, we evaluate the linear model’s ability to reconstruct the cross-encoder scores by examining how well the hypothesized linear model fits the data.
@@ 5.1 Formalizing the Hypothesized Function
In §4.1, we hypothesize that Relevance Scoring Heads appear to compute a summation of Matching Scores weighted by IDF. The Matching Score incorporates soft-TF, along with term saturation and document length signals, all of which are components of the BM25 function. If this hypothesis holds, that these components interact in a BM25-like manner, then the semantic scoring circuit can be expressed as a linear function:
[…]
@@
    $-U_{0}(q_{i})\cdot\text{MS}_{\text{total}}(q_{i},d_{j})$: The interaction term, modeling the product of IDF and TF in BM25.
@@ 5.2 Assessing the Linear Model Fit
The linear regression model achieves a high Pearson correlation (r = 0.8157, p < 0.001) with ground-truth relevance scores, showing that it captures the core of the cross-encoder’s scoring mechanism in a simplified and interpretable form. This correlation surpasses that of the traditional BM25 scoring function under optimized parameters ($k=0.5,b=0.9$; corr = 0.4200, p < 0.001), which demonstrates the discovered $U_{0}$ and $MS$ components effectively capture the signals that the cross-encoder utilizes for ranking. The linear model’s strong generalization to unseen datasets and varying query lengths (details in Appendix §E) further confirms that our circuit understanding captures the cross-encoder’s core mechanism for relevance computation.
@@ 6 Discussion
A two-stage process for relevance computation. Similar to Tenney et al. (2019), who find that BERT rediscovers the classical NLP pipeline, encoding syntax in its lower layers and semantics in its upper, our circuit analysis reveals a corresponding two-stage process. In the early and middle layers, Matching Heads extract lexical signals (e.g., soft-TF matches), reflecting prior work that argues document representations are query-dependent (Qiao et al., 2019). In the upper layers, Contextual Query Representation and Relevance Scoring Heads aggregate these signals into BM25-like relevance scores, consistent with Zhan et al. (2020), who argue that document tokens are largely query-independent. This dual-stage mechanism reconciles these two conflicting viewpoints by showing how query-specific signals migrate from document to query representations across layers and provides a more nuanced understanding of how relevance is computed.
Controllable IR. In §4.4, we show how model editing can be used to upscale or downscale term importance by modifying the encoded IDF values in the embeddings, effectively introducing “tuning dials” for fine-grained control over model behavior. This opens the door to personalization, where models could be adapted to user- or domain-specific preferences without full retraining.
Beyond personalization, these insights enable two additional downstream applications (preliminary experiments in Appendix §F). First, targeted model editing could mitigate unsafe or adversarial content without impairing the model’s ranking effectiveness. Second, aligning internal representations with ground-truth IDF scores could serve as a powerful initialization strategy for fine-tuning, leading to more efficient retrieval pipelines.
@@ 7 Conclusion
Additionally, while the Semantic Scoring circuit bears a strong resemblance to BM25, there may also be other factors influencing relevance computation that are out of the scope of investigation for this work. In this work, we compare it to BM25 because of its well-known status and the fact that previous research has shown that NRMs encode a BM25-like relevance signal (Wang et al., 2021; Rau and Kamps, 2022; Yates et al., 2021; MacAvaney et al., 2022). However, there are other term-matching models (e.g., TF-IDF, QL, etc.) that also rely on similar retrieval heuristics, and it is possible that the model implements these as well.
Furthermore, we note that our BM25-like approximation of the Semantic Scoring circuit does not fully approximate the cross-encoder’s true relevance scores. There are a couple potential explanations for this: (1) the non-linearity of neural models, which our linear regression model does not capture, (2) potentially unexplored components such as additional attention heads that could be uncovered through a different choice of dataset or multi-layer perception (MLP) layers not covered in our analysis, and (3) the incorporation of signals beyond traditional relevance heuristics, such as real-world knowledge learned during pre-training MacAvaney et al. (2022). We leave the investigation of these avenues for future work.
@@ Acknowledgment
Appendix A Detailed Path Patching Methodology
@@
The effect of a patch is measured by the difference in logits (which in the case of cross-encoders, is equivalent to the difference in relevance scores). This algorithm is then iteratively repeated for each important upstream component. For more information, we refer the reader to Wang et al. (2022) and Goldowsky-Dill et al. (2023).
@@
Appendix B Activation Patching on Attention Heads
Appendix C Detailed Path Patching Results
C.1 Path Patching to Final Logits and Intermediate Heads
@@
Figure 8: Activation Patching Results for TFC and STMC diagnostic datasets shows a correlation of 0.96, which suggests that the model employs highly similar components for exact and soft matches.
C.2 Path Patching to Matching Heads
We path patch the residual stream (resid_pre) on each of the Matching Heads using 100 random pairs from the TFC1 and STMC1 datasets (Figures 13, 14, 15, 16). Path patching to the Matching Heads, aimed at tracking the flow of soft-TF information, reveals that at the beginning of the residual stream at layer 0, which corresponds to the embeddings, has the most direct influence on the Matching Heads. In contrast, there is minimal impact from both the attention and MLP layers. This supports the conclusion that the Matching Heads are the primary “generator” of soft-TF, attending to soft matches based on semantic similarity from the embeddings. Thus, we conclude our tracking of soft-TF information, with the Matching Heads identified as the most upstream heads.
Figure 13: TFC1 Path Patching to Matching Heads (8.8, 8.1, 8.0, 7.9, 6.5, 5.9,)
Figure 14: TFC1 Path Patching to Matching Heads (5.7, 4.9, 3.1, 2.1)
Figure 16: STMC1 Path Patching to Matching Heads (5.7, 4.9, 3.1, 2.1)
Appendix D Additional Details on Matching Heads and Additional Signals
Further Patching Experiment to Confirm Matching Heads Are Not Just Writing a Binary Signal of Existence. The signals written by most Matching Heads increase as the number of duplicate tokens rises, rather than remaining constant, proving that these signals are discrete rather than binary. To demonstrate this, we design an activation patching experiment (§4.3) that patches only the Matching Heads using the TFC2 Diagnostic Dataset. Figure 17 shows that for most Matching Heads, increasing the number of duplicate tokens results in a monotonically increasing pattern in their output signals. This behavior further supports our conclusion that Matching Heads compute and encode soft-TF signals rather than merely relaying binary information.
@@ Appendix D Additional Details on Matching Heads and Additional Signals
As the occurrences of $x_{\text{selected}}$ increase, its average attention value grows (Figure 17). On the other hand, the attention for $x_{\text{others}}$ remains relatively constant, which aligns with our soft-TF understanding: only document tokens that are semantically similar to a query token get nonzero attention values, proportional to the extent and frequency of semantic similarity.
@@
Furthermore, for most heads, the TFC2 trend is preserved in the attention values. For LNC1, injecting more irrelevant sentences leads to an increase in the attention value of all query tokens, regardless of whether they match specific query terms (Figure 18). Notably, this increase in attention occurs even as the overall relevance score decreases (as expected from LNC1), suggesting that attention values encode mixed and composite signals of soft-TF, term saturation, and document length, which the model has learned to disentangle to produce appropriate relevance score changes. The varying degrees of influence on the heads’ attention values shown in Figure 17 and 18 suggest that some Matching Heads play a more significant role in regulating these effects, and they may collectively approximate the ideal term effects. Thus, we define this complex attention value as the Matching Score to reflect that it encapsulates soft-TF, term saturation, and document length signals, three important signals of BM25.
Appendix E Generalization across IR datasets
@@
We previously test our linear model on an MSMARCO-based dataset with a fixed number of query tokens as an initial proof of concept. Now, we ask the question: how well does this linear model represent the cross-encoder’s relevance computation on datasets it was not trained on? In this section, we extend our analysis to 12 different IR datasets [^2] Due to computational resources, we use a subset of BEIR datasets Thakur et al. (2021) (ArguAna, Climate-FEVER, FEVER, FiQA-2018, HotpotQA, NFCorpus, NQ, Quora, SCIDOCS, SciFact, TREC-COVID, Touche-2020), but the chosen ones sufficiently cover a wide variety of domains. and investigate various query lengths.
@@
Appendix F SVD Model Editing on Term Importance
In §4.4, we introduce a model editing technique to dynamically control the importance signal of query tokens, and here we present encouraging evidence to use the model editing technique for group-editing the importance signals in downstream tasks.
F.1 Mitigating Adversarial Attacks
Figure 19: Left: Decreasing the IDF of dangerous tokens shows an increasing trend in adversarial mitigation proportion. Right: The intervention maintains a high NDCG score at all levels, demonstrating that the general ranking capability of the cross-encoder remains unaffected.
First, we show the potential of the model editing in mitigating unsafe documents without disrupting the cross-encoder’s ranking performance, unlike retraining or regularization methods that often degrade it.
First, we construct a dataset using obscene and offensive words (LDNOOBW, 2015), filtering out multi-word entries and injecting these unsafe tokens into the safe samples of our TFC datasets. Next, we inspect the subgroup of queries and documents where inserting an unsafe token in the document significantly increases the relevance score. Our goal is to “erase” the effect of the dangerous token by reducing its importance. Among 17,537 adversarial samples, our approach achieves a 80.396% success rate (i.e., the unsafe document is ranked lower than the safe document). For example, when we downweight the target unsafe term by a large factor (e.g., -1200), the model maintains a high nDCG across all ranks, with a score of 0.9861, which is only a 1.39% drop in ranking performance. These results provide preliminary evidence for the potential of localized editing in the IDF-storing low-rank matrix.
F.2 Parameter Efficient Fine-Tuning
Our mechanistic analysis of the model’s relevance mechanisms reveals that term importance (IDF) is encoded in the embedding matrix. Since IDF determines how query terms are weighted in BM25, we hypothesize that only editing or fine-tuning the embedding matrix could provide comparable performance to fine-tuning the entire model, thus providing efficiency benefits.
To test this hypothesis with a preliminary experiment, we use the nfcorpus data, a Nutrition Fact retrieval dataset, from BEIR Thakur et al. (2021). We fine-tune on 10 epochs and evaluate performance across four conditions: (A1) full fine-tuning, (A2) IDF-aligned full fine-tuning (aligning token IDFs to the dataset using the model-editing method from §4.4), (B1) embedding-only fine-tuning, and (B2) IDF-aligned embedding-only fine-tuning. Results across three random (Figure 20 seeds show that IDF-aligned full fine-tuning (A2) achieves the highest NDCG@1, while IDF-aligned embedding-only fine-tuning (B2) outperforms unaligned embedding-only fine-tuning (B1). Additionally, full fine-tuning (A1) and embedding-only fine-tuning (B1) show no significant difference, supporting the idea that adaptation primarily refines the embedding matrix rather than the full model weights. These findings reveal preliminary indications that fine-tuning just on the embedding matrix could potentially be sufficient to adapt cross-encoders to new domains and reveal IDF-aligned initialization as a possible pathway for efficient adaptation.
Appendix G Additional Discussion
Towards a holistic approach in axiomatic analysis. Previous research investigating neural IR models’ adherence to IR axioms shows that BERT does not adhere to most of them, leading to concerns about the limitations of current axiom definitions and the need for new ones Câmara and Hauff (2020). Our findings offer an explanation for these shortcomings, suggesting that the challenge in axiomatic analysis lies with the way it has been applied, rather than the axioms themselves. BERT’s “rediscovery” of BM25 indicates that it leverages a combination of retrieval heuristics to compute relevance, whereas previous analysis treats axioms independently. Going forward, we may consider analyzing axioms in combinations to better reflect the multidimensional nature of relevance.

### v3

@@ 5 Validation of BM25-like Computation
To validate whether Relevance Scoring Heads perform a BM25-style computation as hypothesized in §4.1, we first formalize the hypothesized function of the heads and the information flowing into them as a BM25-style linear function. Next, we evaluate this linear model’s ability to reconstruct the cross-encoder scores by examining how well the hypothesized linear model fits the data.
@@ 5.1 Formalizing the Hypothesized Function
In §4.1, we hypothesize that Relevance Scoring Heads compute a summation of Matching Scores weighted by IDF. The Matching Score incorporates soft-TF, along with term saturation and document length signals, all of which are components of the BM25 function. If this hypothesis that these components interact in a BM25-like manner holds, then the Semantic Scoring Circuit can be expressed as a linear function:
$\sum_{i=1}^{N}$ $\;\text{linear\_combo}\Bigl(-U_{0}(q_{i}),\;\text{MS}_{\mathrm{total}}(q_{i},d_{j}),$ (2)
@@
    $U_{0}(q_{i})\cdot\text{MS}_{\text{total}}(q_{i},d_{j})$: The interaction term, modeling the product of IDF and TF in BM25.
@@ 5.2 Assessing the Linear Model Fit
The linear regression model achieves a high Pearson correlation (r = 0.8157, p < 0.001) with ground-truth relevance scores, showing that it captures the core of the cross-encoder’s scoring mechanism in a simplified and interpretable form. This correlation surpasses that of the traditional BM25 scoring function under optimized parameters ($k=0.5,b=0.9$; corr = 0.4200, p < 0.001), which demonstrates that the discovered $U_{0}$ and $MS$ components effectively capture the signals that the cross-encoder utilizes for ranking. The linear model’s strong generalization to unseen datasets and varying query lengths (details in Appendix G) further confirms that our circuit understanding captures the cross-encoder’s core mechanism for relevance computation.
@@ 6 Discussion
6.1 A Two-Stage Process for Relevance Computation
Similar to tenney2019bert, who find that BERT rediscovers the classical NLP pipeline, encoding syntax in its lower layers and semantics in its upper, our circuit analysis reveals a corresponding two-stage process. In the early and middle layers, Matching Heads extract lexical signals (e.g., soft-TF matches), reflecting prior work that argues document representations are query-dependent (qiao2019understanding). In the upper layers, Contextual Query Representation and Relevance Scoring Heads aggregate these signals into BM25-like relevance scores, consistent with zhan, who argue that document tokens are largely query-independent. This dual-stage mechanism reconciles these two conflicting viewpoints by showing how query-specific signals migrate from document to query representations across layers and provides a more nuanced understanding of how relevance is computed.
6.2 Potential Downstream Applications
In §4.4, we show how model editing can be used to upscale or downscale term importance by modifying the encoded IDF values in the embeddings, effectively introducing “tuning dials” for fine-grained control over model behavior. These insights enable two potential downstream applications.
6.3 Mitigating Adversarial Attacks
Figure 7: Left: Decreasing the IDF of dangerous tokens shows an increasing trend in adversarial mitigation proportion. Right: The intervention maintains a high NDCG at all levels, demonstrating that the general ranking capability of the cross-encoder remains unaffected.
First, targeted model editing could mitigate unsafe or adversarial content without impairing the model’s ranking effectiveness.
As an initial test, we construct a dataset using obscene and offensive words (ldnoobw_badwords), filtering out multi-word entries and injecting these unsafe tokens into the safe samples of our TFC datasets. Next, we inspect the subgroup of queries and documents where inserting an unsafe token in the document significantly increases the relevance score. Our goal is to “erase” the effect of the dangerous token by reducing its importance.
Among 17,537 adversarial samples, our approach achieves a 80.396% success rate (i.e., the unsafe document is ranked lower than the safe document). For example, when we downweight the target unsafe term by a large factor (e.g., -1200), the model maintains a high nDCG across all ranks, with a score of 0.9861, which is only a 1.39% drop in ranking performance.
These results provide preliminary evidence for the potential of localized editing in the IDF-storing low-rank matrix.
6.4 Parameter Efficient Fine-Tuning
Second, aligning internal representations with ground-truth IDF scores could serve as a powerful initialization strategy for fine-tuning, leading to more efficient retrieval pipelines. In other words, rather than updating the entire parameter space, fine-tuning the embedding matrix to better encode IDF values aligned with the retrieval domain may suffice.
To test this hypothesis with a preliminary experiment, we use the nfcorpus data, a Nutrition Fact retrieval dataset, from BEIR beir. We fine-tune on 10 epochs and evaluate performance across four conditions: (A1) full fine-tuning, (A2) IDF-aligned full fine-tuning (aligning token IDFs to the dataset using the model-editing method from §4.4), (B1) embedding-only fine-tuning, and (B2) IDF-aligned embedding-only fine-tuning.
The results in Figure 8 across three random seeds show that IDF-aligned full fine-tuning (A2) achieves the highest NDCG@1, while IDF-aligned embedding-only fine-tuning (B2) outperforms unaligned embedding-only fine-tuning (B1). Additionally, full fine-tuning (A1) and embedding-only fine-tuning (B1) show no significant difference, supporting the idea that adaptation primarily refines the embedding matrix rather than the full model weights.
These findings reveal preliminary indications that fine-tuning just on the embedding matrix could potentially be sufficient to adapt cross-encoders to new domains and reveal IDF-aligned initialization as a possible pathway for efficient adaptation.
@@ 7 Conclusion
Additionally, while the Semantic Scoring circuit bears a strong resemblance to BM25, there may also be other factors influencing relevance computation that are out of the scope of investigation for this work. In this work, we compare it to BM25 because of its well-known status and the fact that previous research has shown that neural IR encode a BM25-like relevance signal (wang2021bert; rau2022different; yates2021pretrained; macavaney2022abnirml). However, there are other term-matching models (e.g., TF-IDF, QL, etc.) that also rely on similar retrieval heuristics, and it is possible that the model implements these as well.
Furthermore, we note that our BM25-like approximation of the Semantic Scoring circuit does not fully approximate the cross-encoder’s true relevance scores. There are a couple potential explanations for this: (1) the non-linearity of neural models, which our linear regression model does not capture, (2) potentially unexplored components such as additional attention heads that could be uncovered through a different choice of dataset or multi-layer perception (MLP) layers not covered in our analysis, and (3) the incorporation of signals beyond traditional relevance heuristics, such as real-world knowledge learned during pre-training macavaney2022abnirml. We leave the investigation of these avenues for future work.
@@ Acknowledgments
Appendix A Activation Patching on Attention Heads
Figure 9: Activation Patching Results for TFC and STMC diagnostic datasets show a correlation of 0.96, which suggests that the model employs highly similar components for exact and soft matches.
Appendix B Example of Semantic Scoring Circuit
Figure 10: Example walkthrough of the hypothesized circuit. We find that the cross-encoder rediscovers a semantic variant of BM25. Specifically, Matching Heads, Contextual Query Representation Heads, and the embedding matrix compute, process and send BM25-like components’ information to the Relevance Scoring Heads which finally compute the relevance score in a BM25-like manner. The trapezoid represents the residual stream representation for each token position.
Figure 10 shows an example of the hypothesized semantic scoring circuit.
In the figure, the model receives an example input: “[CLS] Where is Wellesley [SEP] Wellesley is a college in Massachusetts. [SEP]“.
(1) Matching Heads: These heads generate $MS(\texttt{where})$, $MS(\texttt{is})$, and $MS(\texttt{wellesley})$ that mainly capture query term soft-TF values. For example, the query token wellesley strongly attends to wellesley in the document, and more weakly to related terms like wellesley and college. This attention pattern forms a Matching Score $MS(\texttt{wellesley})$ that also reflects document length normalization (the longer the document, the smaller the soft $MS(\texttt{wellesley})$) and term saturation effects (the more wellesley exists in the document, the less additional gain in $MS$).
(2) Query Contextualization Heads: These heads redistribute the soft-TF signal from key content words to surrounding query tokens. For instance, where and is may receive part of wellesley’s soft-TF signal, allowing the model to amplify high-IDF query terms’ soft-TF signals.
(3) IDF storage: The IDF of all query terms is stored in a dominant low-rank vector of the model’s embedding matrix.
(4) Relevance Scoring Heads: Finally, the model aggregates the reweighted soft-TF signals across query tokens using a mechanism similar to BM25.
Appendix C Detailed Path Patching Methodology
@@
The effect of a patch is measured by the difference in logits (which in the case of cross-encoders, is equivalent to the difference in relevance scores). This algorithm is then iteratively repeated for each important upstream component. For more information, we refer the reader to wang_ioi and goldowskydill_ppatching.
@@
Appendix D Detailed Path Patching Results
D.1 Path Patching to Final Logits and Intermediate Heads
@@
All resulting heatmaps are located at the end of the appendix for clearer flow for the other appendix sections.
D.2 Path Patching to Matching Heads
We path patch the residual stream (resid_pre) on each of the Matching Heads using 100 random pairs from the TFC1 and STMC1 datasets (Figures 19, 20, 21, 22). Path patching to the Matching Heads, aimed at tracking the flow of soft-TF information, reveals that at the beginning of the residual stream at layer 0, which corresponds to the embeddings, has the most direct influence on the Matching Heads. In contrast, there is minimal impact from both the attention and MLP layers. This supports the conclusion that the Matching Heads are the primary “generator” of soft-TF, attending to soft matches based on semantic similarity from the embeddings. Thus, we conclude our tracking of soft-TF information, with the Matching Heads identified as the most upstream heads.
@@ D.2 Path Patching to Matching Heads
Appendix E Additional Details on Matching Heads and Additional Signals
@@
We hypothesize that if the Matching Score contains the term saturation signal, then given more duplicate query terms (TFC2), the summed attention for that term will rise sharply at first and then plateau. As the occurrences of $x_{\text{selected}}$ increase, its average attention value grows (Figure 12). On the other hand, the attention for $x_{\text{others}}$ remains relatively constant, which aligns with our soft-TF understanding: only document tokens that are semantically similar to a query token get nonzero attention values, proportional to the extent and frequency of semantic similarity.
@@
We hypothesize that if the Matching Score contains the term saturation signal, given irrelevant sentences appended (LNC1), the overall attention to the query’s matches will fall. As expected, injecting more irrelevant sentences leads to an increase in the attention value of all query tokens, regardless of whether they match specific query terms (Figure 13). Notably, this increase in attention occurs even as the overall relevance score decreases (as expected from LNC1), suggesting that attention values encode mixed and composite signals of soft-TF, term saturation, and document length, which the model has learned to disentangle to produce appropriate relevance score changes. The varying degrees of influence on the heads’ attention values shown in Figure 12 and 13 suggest that some Matching Heads play a more significant role in regulating these effects, and they may collectively approximate the ideal term effects. Thus, we define this complex attention value as the Matching Score to reflect that it encapsulates soft-TF, term saturation, and document length signals, three important signals of BM25.
Further Patching Experiment to Confirm Matching Heads Are Not Just Writing a Binary Signal of Existence. The signals written by most Matching Heads increase as the number of duplicate tokens rises, rather than remaining constant, suggesting that these signals are discrete rather than binary. To demonstrate this, we design an activation patching experiment (§4.3) that patches only the Matching Heads using the TFC2 Diagnostic Dataset. Figure 12 shows that for most Matching Heads, increasing the number of duplicate tokens results in a monotonically increasing pattern in their output signals. This behavior further supports our conclusion that Matching Heads compute and encode soft-TF signals rather than merely relaying binary information.
Appendix F Additional Details on BM25-like Computation
In this section, we provide additional details on the Relevance Scoring Heads and the Semantic Scoring Hypothesis in §4.1 and §5.
F.1 Correlation Between IDF Values and Attention Distribution
Figure 14: Top: Changes in attention at 10.1 from [CLS] to tok1 when increasing tok1’s IDF at different scales, averaged across all samples. Increasing tok1’s IDF raises attention from [CLS] to tok1, thereby increasing the final score. Bottom: Decreasing tok1’s IDF produces the inverse effect, with slight nonmonotonic deviations suggesting optimal editing windows.
In Section 4.1, we hypothesize that Relevance Scoring Heads combine IDF and soft-TF information to compute final relevance scores in a BM25-like manner. Specifically: (A) increasing or decreasing a token’s IDF leads to (B) the Relevance Scoring Head 10.1222We focus on the attention pattern of 10.1 because it is most positively correlated with IDF (§4.1) and has highest path-patching value. allocating more or less attention from [CLS] to the token, which (C) increases or decreases to the weighting of the token’s soft-TF and thereby the final relevance score.
A $\rightarrow$ B.
[…]
A $\rightarrow$ B $\rightarrow$ C.
In Section 4.4, we modulate IDF by editing the first SVD component of the embedding matrix. Figure 6 shows that scaling a token’s IDF upward or downward produces a monotonic change in the relevance score.
Taken together with Section 5 (which shows that Relevance Scoring Heads perform BM25-style computations on natural IR datasets), these results provide stronger causal evidence for the Semantic Scoring hypothesis.
F.2 Independence between IDF and Soft-TF
In this subsection, we check whether term-frequency (TF) and inverse document frequency (IDF) factors are fully separable.
Low correlation between IDF Soft-TF.
If soft-TF scores are already IDF-weighted, we would expect their one-dimensional values to correlate at least mildly with $U0$. However, across all 13 Matching Heads, the average Pearson correlation between $U0$ and each head’s Matching-Score is low ($r=0.143$, $p<0.01$), suggesting minimal dependence.
Relevance Scoring Heads Do Not Inherit IDF Information From Matching Heads.
Path patching from the Matching Heads to the query vectors of the Relevance Scoring Heads produces almost no change in output score. This indicates that the soft-TF signals produced by the Matching Heads do not strongly convey IDF information, leaving the Relevance Scoring Heads without sufficient signal to prioritize which tokens should receive higher weight.
Appendix G Generalization Across IR Datasets
@@
We previously test our linear model on an MSMARCO-based dataset with a fixed number of query tokens as an initial proof of concept. Now, we ask the question: how well does this linear model represent the cross-encoder’s relevance computation on datasets it was not trained on? In this section, we extend our analysis to 12 different IR datasets [^3] Due to computational resources, we use a subset of BEIR datasets beir (ArguAna, Climate-FEVER, FEVER, FiQA-2018, HotpotQA, NFCorpus, NQ, Quora, SCIDOCS, SciFact, TREC-COVID, Touche-2020), but the chosen ones sufficiently cover a wide variety of domains. and investigate various query lengths.
@@
Appendix H Additional Discussion
Towards a holistic approach in axiomatic analysis. Previous research investigating neural IR models’ adherence to IR axioms shows that BERT does not adhere to most of them, leading to concerns about the limitations of current axiom definitions and the need for new ones camara2020diagnosing. Our findings offer an explanation for these shortcomings, suggesting that the challenge in axiomatic analysis lies with the way it has been applied, rather than the axioms themselves. BERT’s “rediscovery” of BM25 indicates that it leverages a combination of retrieval heuristics to compute relevance, whereas previous analysis treats axioms independently. Going forward, we may consider analyzing axioms in combinations to better reflect the multidimensional nature of relevance.
Figure 19: TFC1 path patching into Matching Heads (layers 8.8, 8.1, 8.0, 7.9, 6.5, 5.9). The residual stream values are patched, with lighter regions indicating stronger path-patching contributions (i.e., where the TFC1 signal flows), and darker regions indicating weaker contributions.
Figure 20: TFC1 Path Patching to Matching Heads (5.7, 4.9, 3.1, 2.1)
Figure 22: STMC1 Path Patching to Matching Heads (5.7, 4.9, 3.1, 2.1)